find the 6th term of the sequence which has a first term of 5 and has a common ratio of 4.2
Answer by RAY100(1637)

Geometric series has form,,,,a(n) = a(1) * r^(n-1)
.
a(1) =5,,,,r=4.2,,,,,n= 6
.
a(6) = 5*4.2^ 5 = 6534.56
.
checking,,5,,21,,,88.2,,370.44,,1555.848,,6534.56,,,,,ok